Solve the system by substitution.
3x – y = –7
x + y = –9

Answer by checkley71(8403)

X+Y=9 OR X=-Y-9
SUBSTITUTE (-Y-9) FOR X IN THE OTHER EQUATION & SOLVE FOR Y.
3(-Y-9)-Y=-7
-3Y-27-Y=-7
-4Y=-7+27
-4Y=20
Y=20/-4
Y=-5 ANSWER.
X-5=-9
X=-9+5
X=-4 ANSWER.
PROOF
3*-4-(-5)=-7
-12+5=-7
-7=-7